a company's profit function, in dollars, is given by P(x)=-0.3x^2+56.16x-638 where x is the number of items sold. in addition, the company has a marginal revenue function of MR(x)=-0.6x+76.16. determine the total cost for the company to produce 10 items.


1
Expert's answer
2021-05-17T08:14:34-0400

The Profit function "P(x)" is the difference between the revenue function "R(x)" and the total cost function "C(x)"


"P(x)=R(x)-C(x)"

Marginal revenue


"MR(x)=R'(x)=-0.6x+76.16"

"R(x)=\int MR(x)dx=\int(-0.6x+76.16)dx"

"=-0.3x^2 +76.16x+k"

"R(0)=0=>k=0"

"R(x)=-0.3x^2+76.16x"

Total cost


"C(x)=R(x)-P(x)"

"C(x)=-0.3x^2+76.16x"

"-(-0.3x^2+56.16x-638)"

"=20x+638"

The total cost for the company to produce 10 items


"C(10)=20(10)+638=838"

The total cost for the company to produce 10 items is "\$838."